40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 14 of 327  Not logged in ELOG logo
ID Date Author Type Category Subject
  15826   Fri Feb 19 16:55:26 2021 KojiUpdateSUSCoM Range on 3"->2" Adapter Ring for SOS

Jordan's screenshot actually shows that the vertical distance (Y) is 0.0000". We want to have the vertical distance of CoM from the wire clamping point to be 0.9mm in the nominal SOS design (this might need to be adjusted to have a similar pitch resonant freq for the different inertia of moment). Let's say it is ~mm ish.

The full range of the bottom dumbbell adjustment gives us the CoM adjustment range of +/-0.002” = +/-50um. This corresponds to an alignment range of +/-50mrad. And we want to set it within +/-500urad.
So we need to adjust the dumbbell position with the precision of 1/100 of the full range (precision of 0.5um).

The groove does not extend to the top of the clamp. The groove shallower than the wire diameter cause the hysteresis of the alignment. Also, the material of the pieces should be stainless steel. Al clamp is softer than the wire and will cause the groove to be dug on the material, causing increased bending friction and hysteresis again.

Saying, all of our suspended masses with Al stand-offs are suffering this issue to some extent. That was the reason to buy the ruby standoffs.

  15825   Fri Feb 19 16:14:16 2021 gautamUpdateSUSCoM Range on 3"->2" Adapter Ring for SOS

I briefly talked with Jordan about this. This suspension will have OSEMs right? With 400ohm series resistance for the coil drivers, we will have ~+/-20mrad actuation range. Of course we'd like to use as much of this for interferometry and not static pitch alignment correction (possibly even increase the series resistance to relax the dewhitening requirements). But what is the target adjustability range in mrad with the dumbell/screw config? My target in the linked elog is 500urad (not any systematic optimum, but will allow us to use most of the DAC range for interferometry). Are these numbers in inches commensurate with this 500urad?

On a related note - are there grooves for the wires to sit in on the side of the sleeve? We looked at the solidworks drawing, and noticed that the groove doesn't extend all the way to the top of the clamp. Also, the material of both the clamping piece and the piece onto which the wire is pressed onto is SS. Don't we want them to be Aluminium (or something softer than the wire) so that the wire makes a groove when the clamp is tightened?

Quote:

We want to move the CoM with the adjustment range so that the residual deviation is adjusted by the bottom dumbbell. 0.0003" is well within the range and good enough.

  15824   Fri Feb 19 16:06:01 2021 KojiUpdateSUSCoM Range on 3"->2" Adapter Ring for SOS

We want to move the CoM with the adjustment range so that the residual deviation is adjusted by the bottom dumbbell. 0.0003" is well within the range and good enough.

 

  15823   Fri Feb 19 15:17:51 2021 JordanUpdateSUSCoM Range on 3"->2" Adapter Ring for SOS

Adjusting the thickness of the cylindrical hole for the mirror on the 2" optic sleeve, from .6875" to .61" thick, moves the CoM to 0.0003" out of plane from the suspension wire. This is with the dumbell at its neutral point.

How close to zero do we need this to be? More fine tuning of that thickness can get it to zero, but this would require much tighter machining tolerance on that hole depth.

Moving the dumbell towards the back of the SOS assembly (noted as negative direction, with origin at the plane formed by the wires), moves the CoM to -0.002" from the plane.

Moving the dumbell towards the front of the SoS assmebly (positive direction wrt the plane formed by the suspension wire), moves the CoM to +0.0022" from the plane.

So the total adjustment range with the dumbell is -0.002"to 0.0022", with the plane formed by the wires as the origin.

See Attachments

Attachment 1: Neutral_Point.png
Neutral_Point.png
Attachment 2: Dumbell_Max_Negative_Travel.png
Dumbell_Max_Negative_Travel.png
Attachment 3: Dumbell_Max_Positive_Travel.png
Dumbell_Max_Positive_Travel.png
  15822   Fri Feb 19 13:38:26 2021 gautamUpdateLSCPRFPMI

I forgot that I had already done some investigation into recovering the PRFPMI lock after my work on the RF source. I don't really have any ideas on how to explain (or more importantly, resolve) the poor seperation of MICH and PRCL sensed in our 3f (but also 1f) photodiodes, see full thread here. Anyone have any ideas? I don't think my analysis (=code) of the sensing matrix can be blamed - in DTT, just looking the spectra of the _ERR_DQ channels for the various photodiodes while a ssingle frequency line is driving the PRM/BS suspension, there is no digital demod phase that decouples the MICH/PRCL peak in any of the REFL port photodiode spectra.

  15821   Fri Feb 19 12:21:04 2021 YehonathanUpdateBHDSOS assembly

A summary of things that need to be fabricated/purchased/done:

Part What needs to be done How much more needed
SUSPENSION BLOCK Fabricate SS dowel pins for 1 suspension block. 2X(diameter 0.094"+-0.002, length 0.38"+-0.01)+2X(diameter 0.188"+-0.002, length 0.5"+-0.01)
WIRE CLAMP If using the opposite side is acceptable, we have enough.  
DUMBBELL STANDOFF Fabricate. Schematics. Need to check the size is compatible with the magnets we have. 40 + 10 for double stacking of side dumbbells. With the existing dumbbells, we'll have 18 spares.
SAFETY STOP, LONG Fabricate or buy. Schematics 4
OSEM assy Check if we have 35. Schematics  
SAFETY STOP, SMALL Fabricate or buy. Schematics 24
SAFETY STOP Fabricate or buy. Schematics 12
SS Spring Plunger Buy from McMaster. Find and check custom plungers around the X arm. 8
4-40 3/8" Ag SHCS Buy from uccomponents.com 30
4-40 1/2 Ag SHCS Buy from uccomponents.com 60
1/4-20 3/4 Ag SHCS Buy from uccomponents.com 150
1/4-20 5/4 Ag SHCS Buy from uccomponents.com 30
1/4 SS Lock Washer Buy from McMaster 30
1/4 SS Lock Wassher (Reduced OD) Buy from McMaster 30
Viton Tips Need to find stock Not sure. Existing eq stops have phosphor bronze springs. Should all of them be replaced with Viton?
Steel Music Wire There are 500ft of wire (enough for many SOSs) in a desiccator somewhere according to this elog  

 

  15820   Thu Feb 18 20:24:48 2021 KojiSummaryElectronicsA bunch of electronics received

Todd provided us a bunch of electronics. I went to Downs to pick them up this afternoon and checked the contents in the box. Basically, the boxes are pretty comprehensive to produce the following chassis

  • 8 HAM-A coil driver chassis
  • 7 16bit Anti-Aliasing chassis
  • 4 18bit Anti-Imaging chassis
  • 5 16bit Anti-Imaging chassis

Some panels are missing (we cannibalized them for the WFS electronics). Otherwise, it seems that we will be able to assemble these chassis listed.
They have placed inside the lab as seen in the attached photo.


HAM-A COIL DRIVER (Req Qty 28+8)

- 8 Chassis
- 8 Front Panels
- 8 Rear Panels
- 8 HAM-A Driver PCBs
- 8 D1000217 DC Power board
- 8 D1000217 DC Power board

16bit AA (Req Qty 7)
- 7 CHASSIS
- 6 7 Front Panels (1 missing -> [Ed 2/22/2021] Asked Chub to order -> Received on 3/5/2021)
- 7 Rear Panels
- 28 AA/AI board S2001472-486, 499-511
- 7 D070100 ADC AA I/F
- 7 D1000217 DC Power board

18bit AI (Req Qty 4)
- 4 CHASSIS
- 4 Front Panels
- 4 Rear Panels
- 8 AA/AI board S2001463-67, 90-92
- 4 D1000551 18bit DAC AI I/F
- 4 D1000217 DC Power board
- bunch of excess components

16bit AI (Req Qty 5)
- 5 CHASSIS
- 4 5 Front Panels (D1101522) (1 missing -> [Ed 2/22/2021] Asked Chub to order -> Received on 3/5/2021)
- 3 5 Rear Panels (D0902784) (2 missing -> [Ed 2/22/2021] Asked Chub to order -> Received on 3/5/2021)
- 10 AA/AI board S2001468-71, 93-98
- 5 D1000217 DC Power board
- 5 D070101 DAC AI I/F

Internal Wiring Kit

[Ed 2/22/2021]
Asked Chub to order:
- Qty 12 1U Hamilton Chassis
- Qty 5 x Front/Rear Panels/Internal PCBs for D1002593 BIO I/F (The parts and connectors to be ordered separately)

  -> Front/Rear Panels received (3/5/2021)
  -> PCBs (unpopulated) received (3/5/2021)
  -> Components ordered by KA (3/7/2021)

Attachment 1: IMG_0416.jpeg
IMG_0416.jpeg
  15819   Thu Feb 18 20:20:25 2021 KojiUpdateSUSaLIGO Sat Amp characterization

Yeah, it's really inconsistent. You had 35mA LED drive and the current noise of the noisy channel was 5e-7 A/rtHz at 1Hz. The RIN is 1.4e-5 /rtHz. The approx. received photocurrent is 30uA as we discussed today and this should make the noise around 4e-10 A/rtHz at 1Hz. However, the readout noise level is better than this level. (well below 1e-10 A/rtHz)

BTW, the IMC seemed continuously locked for 5 hours. Good sign.

  15818   Thu Feb 18 18:05:04 2021 gautamUpdateSUSaLIGO Sat Amp characterization

Before installation, I performed a bunch of tests on the aLIGO sat amp. All the measurements were made with the dummy suspension box substituting for an actual suspension. Here are the results.

Attachment #1: Transimpedance amplifier noises.

  • Measurement setup: J7 of the Satellite Amp goes to J9 on D1900068 front end (even though the connector is actually labelled "J3" on the box we have - maybe a versioning problem?). The outputs then go to a G=100 SR560 in AC coupled mode (the main purpose here was to block the large DC from the SR785, but I tacked on G=100 while I was at it).  
  • Top panel shows the raw measured voltages.
  • The bottom panel does a bunch of transformations:
    • Undoes the z:p = 3:30 Hz whitening on board the sat amp.
    • Undoes the G=100 gain of the SR560, and the AC coupling poles/zeros of SR560 and SR785.
    • Converts from voltage to current by dividing by the transimpedance gain, 242 kohms. 
  • Some model curves are shown for comparing to the measured spectra. It may be possible that we don't need to modify the nominal z:p = 0.4:10 Hz - I don't think the nominal seismic level will saturate the output even with the 0.4:10 Hz whitening, and it gives us even more clearance to the ADC noise (although we don't need it, we are gain limited at those frequencies, this is mostly a suggestion to reduce the workload).
  • The neon green curve is measured with the actual MC1 suspension plugged in, local damping enabled. It doesn't line up with the nosie floor of the bench tests, probably because the cts/um conversion factor could be off by some factor? Around 1 kHz, you can also see some broad peaks that are reminiscent of those seen in the MC_F spectrum after the c1psl Acromag upgrade. I hypothesize this is due to some poor grounding. Hopefully, once we get rid of the single-ended sending/receiving components in the suspension electronics chain, these will no longer be an issue.

Attachment #2: LED drive current source noises. I mainly wanted to check a claim by Rich in a meeting some time ago that the LED intensity fluctuations are dominated by inherent LED RIN, and not by RIN on the drive current. 

  • Measurement setup: a pair of pomona mini-grabbers was used to clip onto TP3. I found the voltage noise to be sufficiently high that no preamplification was required, and the DC level was <1V, so I just used the SR785 in AC coupled mode. 
  • The dummy suspension box was being driven while the measurement was being made (so the current source is loaded).
  • One channel (CH6) shows anomalously high nosie. I confirmed this was present even after the box was plugged in for ~1 day, so can't be due to any thermal / equilibriating transients.
  • I didn't check for consistency at the monitor testpoint, but that is exposed even with the MC1 suspension plugged in, so we can readily check. Anyways, from the corresponding photodiode curve in Attachment #1, it would seem that this excess RIN in the drive current has no measurable effect on the intensity fluctuations of the LED (the DC value of the paired PD is consistent with the others, ~6V DC). I must say I am surprised by this conclusion. I also checked for coherence between TP3 and the PD output using the SR785, and found none. 🤔 
  • Nevertheless, for the remaining channels, it is clear that the drive current is not shot noise limited for <1kHz. This isn't great. One possible reason is that the collector voltage to Q1 is unregulated (my modeling suggests only ~10dB rejection of collector voltage fluctuations at the output). I believe the current source designed by Luis for A+ makes some of these improvements and so maybe Rich was referring to that design, and not the aLIGO Satellite Amplifier flavor we are using. Anyways, this is just academic I think, the performance is the unit is fine for our purposes.

I will update with the MC1 suspension characterization (loop TFs, step responses etc) later.

Attachment 1: OSEMnoise.pdf
OSEMnoise.pdf
Attachment 2: LEDdriveNoise.pdf
LEDdriveNoise.pdf
  15817   Thu Feb 18 15:33:21 2021 gautamUpdateSUSaLIGO Sat Amp installed, powered and commissioned

The WFS servo was recommissioned. The matrix can be tuned a bit more, but for now, I've recovered the old performance and the alignment doesn't seem to be running away, so I defer further tuning for later. The old Satellite box was handed over to Yehonathan for his characterization of the "spare" OSEMs.

This finishes the recovery of the MC1 suspension, I am now satisfied that the local damping loops are performing satisfactorily, that the WFS servo is also stable, and that POX/POY locking is recovered. On MC1, we even have 4 actuatable face OSEMs and the PIT(YAW) bias adjust slider even moves the optic in PIT(YAW), what a luxury. 

I've SDFed all the changes, and have backup of the old realtime model and C1SUSAUX_MC1 database files if we want to go back for whatever reason. The changes required to make this suspension work are different from what will eventually be required for the BHD suspensions (because of the hybrid iLIGO/aLIGO electronics situation), so I will not burden the readers with the tedious details.

  15816   Thu Feb 18 15:15:12 2021 yehonathanUpdateSUSOSEM testing for SOSs

I am setting up a testing rig for the OSEMs we recently obtained. I found the schematic for the OSEM assembly from which the pin assignment can be read.

I connected the OSEM's pin plate to a female DB15 on a breakout board. I find the pin assignment (attachment 1, sorry for the image quality) to be:

1 PD Cathode
2 LED Anode
3 Coil end
4 PD Anode
5 LED Cathode
6 Coil Start

There are several things that need to be done for each OSEM.

1. Measuring inductance of the coils. I checked that the measurement wires don't add any measurable inductance.

2. Check that the PDs and LEDs are alive (e.g. check forward voltage drop with fluke)

3. Energize the LED and PD.

4. Check PD DC level. For this, I might need the satellite box amplifier.

5. Check LED spot position on the PD.

6. Re-engrave OSEM S/N if needed.

OSEM # Coil Inductance (mH) Coil resistance (ohm) PD forward voltage (V) LED forward voltage (V)
280 2.87 14.1 0.63 1.1
         
         

I still need to figure a sensible scheme for points 3-5.

 

 

Attachment 1: OSEM_Pin_Plate.png
OSEM_Pin_Plate.png
  15815   Thu Feb 18 03:20:09 2021 KojiSummaryElectronicsCurrent Rack Map

For your planning:

Attachment 1: rack_plan.pdf
rack_plan.pdf
  15814   Wed Feb 17 16:11:53 2021 gautamUpdateSUSaLIGO Sat Amp installed, powered

There is some non-trivial sign flipping in the sensors/coils in this new setup because it is a hybrid one with the old interfacing electronics (D000210, D010001) and the new Satellite Amplifier (D080276). So I haven't yet gotten the damping working. I am leaving the PSL shutter closed and will keep working on this today/tomorrow. I have made various changes to the c1mcs realtime model and the c1susaux database record where MC1 is concerned. I have backups of the old ones so we can always go back to that if we so desire.

In the meantime, the PSL shutter is closed and there is no light to the IFO.


Update 1700: I've implemented some basic damping and now the IMC is now locked. The WFS loop runs away when I enable it, probably some kind of weird interaction with the (as of now untuned) MC1 local damping loops. I will write up a more detailed report later.


Update 2300: Did the following:

  1. Re-calibrated the cts2um filter in all SENSOR filter banks to account for the increased transimpedance and LED drive current. I judged the overall scaling to be x0.25 but this can be calibrated against the bounce peak height for example (it lines up pretty well).
  2. Re-measured the input matrix - it was very different from what was loaded. I am measuring this again overnight for some consistency.
  3. Re-tuned the damping gains. Now the optic damps well, and the loops seem file to me, both via broadband noise injection TF and by step response metrics.
  4. Yet, the WFS servo cannot be enabled. The WFS signal is summed in before the output matrix so I don't know why this would have a different behavior compared to the local damping, or indeed, why this has to be changed. Will need some (WFS) sensing/actuation matrix measurements to know more.

Dropping this for tonight, I'll continue tomorrow. Meanwhile, the OSEM input matrix measurement is being repeated overnight. PSL shutter is closed.

  15813   Wed Feb 17 13:59:43 2021 KojiUpdateSUSCoM on 3"->2" Adapter Ring for SOS

Note from today's meeting:

1. Can we adjust the thickness of the cylindrical hole for the mirror to move the COM in the plane of the wires. (We should be able to do that)

and

2. Please check how much we can displace the COM by the bottom dumbbell.

  15812   Wed Feb 17 13:59:35 2021 gautamUpdateDetCharSummary pages

The summary pages had failed because of a conda env change. We are dependent on detchar's conda environment setup to run the scripts on the cluster. However, for some reason, when they upgraded to python3.9, they removed the python3.7 env, which was the cause of the original failure of the summary pages a couple of weeks ago. Here is a list of thoughts on how the pipeline can be made better.

  1. The status checking is pretty hacky at the moment.
    • I recommend not using shell via python to check if any condor jobs are "held".
    • Better is to use the dedicated python bindings. I have used this to plot the job durations, and it has worked well.
    • One caveat is that sometimes, there is a long delay between making a request via a python command, and condor actually returning the status. So you may have to experiment with execution times and build some try/except logic to catch "failures" that are just the condor command timing out and not an actual failure of the summare jobs.
  2. The status check should also add a mailer which emails the 40m list when the job is held. 
    • With htcondor and python, I think it's easy to also get the "hold reason" for the job and add that to the mailer.
  3. The job execution time command is not executing correctly anymore - for whatever reason, the condor_history command can't seem to apply the constraint of finding only jobs run by "40m", although running it without the constraint reveals that these certainly exist. Probably has to do with some recent upgrade of condor version or something. This should be fixed.
  4. We should clear the archive files regularly. 
    • The 40m home directory on the cluster was getting full. 
    • The summary page jobs generate a .h5 archive of all the data used to generate the plots. Over ~1 year, this amounts to ~1TB.
    • I added the cleanArchive job to the crontab, but it should be checked.
    • Do we even need these archives beyond 1 day? I think they make the plotting faster by saving data already downloaded locally, but maybe we should have the cron delete all archive 
  5. Can we make our own copy of the conda env and not be dependent on detchar conda env? The downside is that if something dramatic changes in gwsumm, we are responsible for debugging ourselves.

Remember that all the files are to be edited on nodus and not on the cluster.

  15811   Tue Feb 16 22:59:36 2021 YehonathanUpdateBHDSOS assembly

Done.

Also, the magnets are nickel-plated. I guess that doesn't matter for the baking (Curie temp of 355 °C)?

Quote:

The curie temp of SmCo seems about x2 (in K) of the one for NdFeB. i.e. 600K vs 1000K. So I believe 177degC = 450K is not an issue. Just make sure the curie temp, referring the specific property for the magnets from this company. (You already know the company from the procurement doc). It'd be great if you upload the doc on the 40m wiki.

 

  15810   Tue Feb 16 15:29:01 2021 KojiUpdateBHDSOS assembly

The curie temp of SmCo seems about x2 (in K) of the one for NdFeB. i.e. 600K vs 1000K. So I believe 177degC = 450K is not an issue. Just make sure the curie temp, referring the specific property for the magnets from this company. (You already know the company from the procurement doc). It'd be great if you upload the doc on the 40m wiki.

  15809   Tue Feb 16 14:56:44 2021 gautamUpdateSUSaLIGO Sat Amp installed, powered

[jordan, gautam]

  • Ran 60ft long cables from 1X4 to MC1/MC3 chamber flange, via overhead cable tray, and top of PSL enclosure for the last ~20ft. Note that it may be that the overhead cable trays cannot support the weight of the cables for 15 SOSs (total 30 shielded cables with 37 wires as twisted pairs) when we eventually add the optics for the BHD upgrade.
  • Installed aLIGO satellite amplifer in 1X4.
  • Tapped +/-20 V (which is the available voltage closest to the required +/-18V). For this, the Sorensens were powered down, and the actual taps were made from the fusable blocks powering the Trillium interface box. We made sure to leave an extra slot so that this kind of additional headache is not required for the next person doing such work.
  • Once installed, I plugged in the dummy suspension box and verified that the unit performs as expected. 
  • Some photos of the installation are here.

After this work, the IMC locked fine, the AS camera has the Michelson fringing, the fast CDS indicators are all green, and the seismometer BLRMS all look good - therefore, I claim no lasting damage was done as a direct result of today's work at 1X4. I will connect up the actual suspension at my leisure later today. Note that the MC1 glitches seem to have gone away, without me doing anything about it. Nevertheless, I think it's about time that we start testing the new hardware. 


Unrelated to this work: while I was testing some characteristics of the MC1 suspension (before we did any work in the VEA, you can see the timestamp in the ndscope), I noticed that the MC1 UL coil channel cannot actually be used to actuate on the optic. The coil driver Vmon channel demonstrates the appropriate response, which means that the problem is either with the Satellite box (it is just a feedthrough, so PCB trace damaged?) or with the OSEM itself (more likely IMO, will know more once I connect the new Satellite Amplifier up). I only show comparison for UL vs UR, but I checked that the other coils seem to be able to actuate the optic. This means we have been running for an indeterminate amount of time with only 3 face actuators on MC1, probably related to me having to do this work


Also unrelated to this work - while poking around at 1X5 rear, I noticed that the power connections to the existing Satellite Boxes are (understatedly) flaky, see connections to T1-T4 in Attachment #2..

Attachment 1: MC1_deadUL.png
MC1_deadUL.png
Attachment 2: IMG_9100.jpg
IMG_9100.jpg
  15808   Tue Feb 16 13:13:39 2021 YehonathanUpdateBHDSOS assembly

Gautam pointed out that there are extra Sm-Co magnets stored in the clean optics cabinet.

I took the magnet box out and put it on the rolling table next to south flow bench. The box contains 3 envelopes with magnets.

They are labelled as following:

FLUX 94 - 50 parts

FLUX 93 - 10 parts

FLUX 95 - 40 parts

(What is FLUX??)

The box also contains some procurement documents.

The clean and bake dcc says :

1. Ultrasonic clean in methanol for 10 minutes

2. Bake in vacuum at 177 C° for 96 hours

Should we go ahead with the C&B?

  15806   Fri Feb 12 15:03:48 2021 JordanUpdateSUSCoM on 3"->2" Adapter Ring for SOS

As it currently stands the Center of Mass of the Adapter Ring/Optic assembly is 0.0175" out of the plane formed by the suspension wire. See Attachments. The side plate, along with the EQ stops are hidden to show the CoM and the plane.

Note: The changes discussed in the meeting with Calum have not been added and are a work in progress. These changes include:

- Adding a 45 deg chamfer to the both parallel faces of the adapter ring. This along with a modified bracket for the EQ stops will allow for easier adjustment of the screws. 

- Potentially changing material of adapter ring to stainless stell to more accurately emulate the mass of a 3" optic.

- Different adjustment mechanism of the "dumbell" at bottom of adapter ring to something similar to the VOPO suspension (will need to consult Calum further)

Attachment 1: Screenshot_(1).png
Screenshot_(1).png
Attachment 2: Screenshot_(3).png
Screenshot_(3).png
Attachment 3: CoM.PNG
CoM.PNG
  15805   Thu Feb 11 18:21:39 2021 gautamUpdateSUSMC suspension glitches

MC1 suspension is glitching again, so this is a good chance to install the new sat box and test it in the field.

  15804   Thu Feb 11 16:58:52 2021 ranaSummaryBHDSatellite Amplifier Very Low frequency noise After modifications

I expect that a single OSEM channel can't be better than 1e-10 m/rHz above 5 Hz, so probably something wrong in the calibration. 1.6 V/mm seems right to me, so could be some place else.

  15803   Thu Feb 11 11:10:05 2021 AnchalSummaryBHDSatellite Amplifier Very Low frequency noise After modifications

Here is a proper measurement for PD transimpedance amplifier circuit in the Satellite amplifier box D1002818 S2100029. The input from rear DB25 connector was left open and measurement was taken with AC coupling with correction by the AC coupling transfer function (Zero at 0, pole at 160 mHz). I have calculated the input referred displacement noise by calculating the conversion factor of OSEM in A/m. From 40m/12470, old conversion factor of OSEM to output of sat amplifier was 1.6 V/mm. then, the transimpedance was 39.2 kOhm, so that must mean a conversion factor of 1.6e3/39.2 A/m. This I scaled with increased drive current by factor of 35/25 as mentioned in this document. The final conversion factor turned out to be around 57 mA / m. If someone finds error in this, please let me know.

There is excess noise in the low-frequency region below 5-6 Hz. If people think I should make a measurement of amplified noise to go further away from the instrument noise floor, let me know.

Attachment 1: AfterChangesSpectrum_AC.zip
Attachment 2: D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf
D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf
Attachment 3: D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf
D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf
Attachment 4: D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf
D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf
  15802   Wed Feb 10 21:14:03 2021 gautamUpdateElectronicsProduction version of the HV coil driver tested

Summary:

I did what I consider to be a comprehensive set of tests on the production version of the high voltage coil driver circuit. I think the performance is now satisfactory and the circuit is ready for the production build. Barring objections from anyone, I will ask Chub to place an order for components to stuff the 4 necessary units + 1 spare on Friday, 12 Feb (so that people have a full day to comment). A big thanks to Chub and the folks at JLCPCB for dealing with my incessant order requests and patiently supporting this build and letting me turn this around in 10 days - hopefully this is the end of this particular saga.

Schematic is here. All references to component designations are for v4 of the schematic.

Important design changes:

  1. All I/O to this board will be via D9 connectors. This will allow bypassing the high voltage stage in future suspensions while retaining the same cable config in the suspension drive, if that is desired. Some re-arrangement of the grouping of coils was also done for consistency with the planned upgrade.
  2. Differential receiving for the input signal from the Acromag. The nominal quad opamp is LT1125 but if we find noise issues (which I didn't), the OP497 has compatible PCB footprint.
  3. Added input protection dual diode D6 to protect the PA95 as recommended in the datasheet. This should protect the IC if (for example) the HV line isn't plugged in but the Acromag input is non-zero.
  4. Increased the feedback resistance from 30kohms to 12kohms. R16 through R21 are now 20k, while the old design had 5k. The purpose is to reduce the current demand in the feedback path, hopefully this opens up the number of DCPS we can use. To keep the overall gain of 31, the resistor R15 was upped from 1kohms to 4kohms.
  5. Feedback capacitance reduced from 15 uF to 3 uF. This was largely for space saving / ease of layout on the PCB. The resulting corner frequency is increased slightly from 0.35 Hz to 0.45 Hz but this doesn't have any imapct on the performance of the circuit at frequencies of interest (1/2/pi/R/C had R=30k, C=15uF, now R=120k, C=3uF).
  6. Added an R-C-R network at the output of the PA95, before the fast actuation signal is summed and sent to the OSEM coil.
    • This is probably the most important change, noise-performance wise.
    • The purpose of the network is to passively filter out the excess noise we saw at ~100 Hz (the corner from the 4kohm resistor + 3uF cap is at ~13 Hz, so factor of 10 filtering at 100 Hz, which was deemed sufficient, see earlier elogs in the thread). 
    • The Johnson noise contribution of the 20 kohm resistor is slightly higher than the original design which had a 25 kohm series resistor (but no R-C-R passive filter at the output of the PA95). But once again, this was deemed to have negligible effect on the performance at frequencies of interest to us.
    • The total current driving capability of the circuit is almost unchanged since the 20kohm + 4kohm nearly equals the old 25kohm resistance.
  7. Made the Vmon paths for monitoring the HV output of the PA95 differential sending, seems like a good practise to follow for all new designs.
  8. Added on-board bypass capacitors (2 x 10uF WIMA film caps) for cleaning up the HV supply noise.

Tests:

A series of tests were done. Note that only 1 channel was stuffed (I am out of PA95s), and the HP power supplies borrowed from Rich were used for the HV rails. For the +/-18V, a regular bench-top unit was used.

  1. Low voltage stage tests
    • TF of the differential receiving stage was measured and the overall unity gain and corner at 24kHz were verified, see Attachment #1.
    • With the input of the circuit grounded, I measured the noise of the circuit at various points (see legends on Attachment #2). I didn't bother to do a detailed verification against a SPICE model as the levels seemed roughly what is expected.
  2. Overall noise performance with HV stage enabled
    • For a range of drive currents, generated by applying the appropriate voltage using an Acromag XT1541 DAC module to the J1 connector, I measured the voltage on the circuit side of the 20 kohm resistor (by clipping onto the resistor leg. Note that the path to ground for the current was provided by connecting a 20 ohm resistor between pins 1 and 6 on J3a, and then grounding pin 6.
    • Results are shown in Attachment #3
    • For the drive currents at the edge of the range of operation, there is a small excess relative to lower drive currents. My best hypothesis for why this is happening is that the HV rail is too close to the requested output voltage (the HP units are rated for 320V, which is borderline if we want 300V at the output of the PA95). In any case, the R-C-R passive filter means that above ~60 Hz, there is excellent agreement between model and measurement.
  3. Time domain tests
    • Used a function generator. to drive a 50 mHz, 3Vpp sine wave to the "Bias Input" (=J1), and monitored (i) pickoff of drive signal, (ii) High voltage output at the circuit side of the 20kohm resistor, and (iii) the Vmon output (=pins 1/6 on J4), all on a 100 MHz Tektronix scope.
    • Results shown in Attachment #4. Once again, I see no red flags.
    • While I had the unit hooked up to the scope, I also checked the time domain signal with the scope set to 100 ns/div time base. I saw no evidence of any oscillatory features, either when no input signal was applied, or when a DC signal was provided (in which case the scope was set to AC coupling). 👍 
  4. Checked that the protection diodes at various locations in the circuit work.
  5. Checked the pin-mapping on all 6 D9 connectors is consistent with the top level diagram in the schematic.

PCB remarks:

As I was stuffing the board, I noticed a few improvements that can be made. Just noting these here for documentation purposes - these changes are mostly aesthetic and I personally see no need to order another set of PCBs.

  1. In some places, the silkscreen designators don't have the correct "orientation" relative to the component they are designating. I didn't find any serious ambiguity in terms of being misled to stuff the wrong component onto the wrong pads, but in the spirit of doing a professional job...
  2. The current limiting resistors on the +/-430V LEDs (R37/R38) have footprints for leaded components rather than SMT (which is what the +/-15V LEDs have).
  3. R45 and R46, the current limiting resistors for the rear panel power indicator LEDs, have 0805 footprint rather than 1206.
  4. When I drew up the PCB, R23, the 4kohm resistor in the R-C-R network, was set up as a 1W resistor. Let's say there can be 15 mA flowing through this resistor - the power dissipated is 15e-3 ^2 * 4e3 is 0.9W, which is uncomfortably close to the limit. For all the tests above, I used a 3W resistor, and didn't find any serious noise issues. The drilled holes are a little tight for this higher power rated resistor, but I don't think this is a showstopper.

Communications with Apex:

I've been talking to support at Apex, and pointed out that I couldn't match the SPICE model performance even for a simple non-inverting amplifier with the PA95. The feedback I got from them was that 

  1. They don't optimize the SPICE models for input noise and so it was a nice coincidence that model and measurement are somewhat close (but not exactly).
  2. They recommend the PA194, which is actually advertised as "low-noise". The PA95 is apparently not a "low-noise" part, with its 2uVrms input noise. 

Whiel the PA194 is compatible with our voltage and current requirements for this application, it is ~3x the cost, and seems like the R-C-R output filter allows us to realize the goal of 1pA/rtHz, so I'm inclined to stick with the PA95.

Production assembly:

I'd prefer to get as much of the board stuffed by Screaming Circuits as possible. It took me ~3 hours to stuff 1 channel + the power supply parts, standoffs etc. So I estimate it'll take me ~6 hours to stuff the entire board. So not the end of the world if we have to do it in-house.

Attachment 1: inputDiffRecTF.pdf
inputDiffRecTF.pdf
Attachment 2: LVnoises.pdf
LVnoises.pdf
Attachment 3: totalNoise.pdf
totalNoise.pdf
Attachment 4: timeDomainTests.pdf
timeDomainTests.pdf
  15801   Wed Feb 10 17:18:03 2021 KojiSummaryBHDSatellite Amplifier Output Offset measurements

Testing the satellite amp i.e.  PD driver
- To test the noise of the PD transimpedance amps:
Leave the PD input open (do not short the terminal goes to the PD)
- To test the current noise of the LED drivers: Short the output with an appropriate Rs to have the nominal current.
- To test the overall noise level together with the LED/PD pair: Connect the dummy OSEM module.

Testing the coil drivers
-
Short the output with an appropriate Rs.

  15800   Wed Feb 10 15:25:45 2021 gautamSummaryBHDSatellite Amplifier Output Offset measurements

Why not just do this test with the dummy suspension box and CDS system? I think Rich's claim was that the intrinsic LED RIN was dominant over any drive current noise but we can at least measure the quadrature sum of the two (which is after all the relevant quantity in terms of what performance we can realize) and compare to a model.

  15799   Wed Feb 10 15:07:50 2021 AnchalSummaryBHDSatellite Amplifier Output Offset measurements

I measured the output DC voltage of the satellite amplifier box at PDMon port when the PDA input was shorted and got following offsets:
 

CH Output Offset (mV) CH Output Offset (mV)
1 6 5 750
2 140 6 120
3 350 7 537
4 40 8 670

However, I think I'm making a mistake while measuring this offset as well as all the noise measurements of this satellite amplifier box so far. Since it is a current input, transimpedance circuit, the noise of the circuit should be measured with open input, not closed. Infact, by shorting the PDA input, I'm giving DC path to input bias current of AD833 transimpedance amplifier to create this huge DC offset. This won't be the case when a photodiode is connected at the input which is a capacitor and hence no DC path is allowed. So my issue of offset was bogus and past two noise measurements in 40m/15797 and 40m/15793 are wrong.

  15798   Wed Feb 10 14:14:58 2021 gautamUpdateElectronicsCustom cables received

We received the custom cables to test the new suspension electronics. They are under my desk. So we are ready.

This batch was a small one - the company says that they can make molded cables if we have a minimum order, something to consider I gues.s.


Update 1900 11 Feb: I verified that the pin outs of the cables are as we intended (for one set of each type of cable). Because this was a small order, the connectors have metal shells, and so for cable #2 (sat box to flange), the two shells are shorted to each other. I can't verify if the shield is isolated from the shell on J5 without cutting open the cable. One thing that occurred to me is that we should give pins 5,8,11 on J4 and 16,20,24 on J5 (respectively) unique identifiers. They should only be shorted to GND on the circuit board itself. To be fixed for the next iteration. I uploaded some photos here.

I was unable to measure the capacitance of the cable using the LCR meter, and didn't opt to try any other method.

Attachment 1: satWiring.pdf
satWiring.pdf
  15797   Wed Feb 10 11:45:59 2021 AnchalSummaryBHDSatellite Amplifier Very Low frequency noise After modifications

As suggested, I wrapped the satellite amplifier box D10028128 S2100029 in blanket and foam and took very low frequency spectrum starting from 32 mHz to 3 Hz. The results are attached along with stiched high frequency measurements from 40m/15793.

Very Low Frequency Spectrum Measurement

  • D1002818 S2100029 box was powered and covered in a foam blanket.
  • Additionally, it was covered from all sides with foam to reduce wind and temperature effects on it.
  • The rear panel DB25 connector was connected to a breakout board where pins od PDA input and GND were shorted, shorting the transimpedance circuit input.
  • The output was read from PDMon DB9 output at front panel which was converted to 4 BNC channels using breakout board.
  • Two channel noise was measured at once using D1002818_SP.yml parameter file.
  • Instrument noise at all the used input ranges were measured separately by shorting the input of the BNC cables.

Edit Wed Feb 10 15:14:13 2021 :

THIS MEASUREMENT WAS WRONG. SEE 40m/15799.

Attachment 1: FrontsideLook.jpg
FrontsideLook.jpg
Attachment 2: BacksideLook.jpg
BacksideLook.jpg
Attachment 3: InnerFoamBlanket.jpg
InnerFoamBlanket.jpg
Attachment 4: D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf
D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf
Attachment 5: D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf
D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf
Attachment 6: AfterChangesLFSpectrum.zip
  15796   Thu Feb 4 15:14:55 2021 YehonathanUpdateBHDSOS assembly

I gathered all the components I could find from the SOS towers and the cleanroom and put it all on the table next to the flow bench (See attachment).

I combed through the cleanroom cabinet for SOS parts but didn't find all the parts listed in the procurement spreadsheet. I did find some extra items that were not listed.

This table compares the quantities in the spreadsheet to the quantities collected on the table. Green rows are items I found more than in the procurement spreedsheet while red rows are items I found less.

ITEM DCC # Qty required Qty in procurement spreadsheet How much I found
SENSOR/ACTUATOR PLATE D960002 14 21 21
SUSPENSION BLOCK D960003 7 9 9
TOWER BASE D960004 7 10 11
RIGHT SIDE PLATE D960005 7 12 13
LEFT SIDE PLATE D960006 7 12 13
UPPER MIRROR CLAMP D960007 7 8
7+1 teflon
LOWER CLAMP D960008-1 7 8 8
LOWER CLAMP, OPPOSITE D960008-2 7 8 8
WIRE CLAMP 1205308-1 10 17 9
CLAMP, SUSPENSION BLOCK D960134 14 19 21
STIFFENER PLATE D960009 7 9 9
DUMBBELL STANDOFF D970075 50 10 7
SAFETY STOP, LONG D970313 14 2 10
OSEM assy D960011 35 2 13 wire wound osem housings (gold)
WIRE STANDOFF D970187 20 7 0
GUIDE ROD D970188 10 9 0
MAGNET D960501 50 54 51 rusted + 37 slightly rusted. Didn't put on table
SAFETY STOP, SMALL D970312 28 0 4
SAFETY STOP D970311 28 0
16+9 stained w/o spring
SS Spring Plunger 8498A999 35 4 27
Attachment 1: 20210204_144007.jpg
20210204_144007.jpg
  15795   Wed Feb 3 21:28:02 2021 gautamUpdateCDSCDS crash and CDS/IFO recovery

I am just reporting my experience - this may be a new failure mode but I don't think so. In the new RTCDS, the ntp server for the FEs are the FB, to which they are synced by timesyncd. The FB machine itself has the ntpd service installed, and so is capable of synching to an NTP server over the internet, but also serving as an NTP server for the FEs. The timesyncd daemon may not have started correctly, or the NTP serving from FB got interrupted (for example), but that's all just speculation.

  15794   Wed Feb 3 18:53:31 2021 KojiUpdateCDSCDS crash and CDS/IFO recovery

Really!? I didn't reboot the machines between "sudo date" and "rtcds start c1x0*". I tried rtcds. If it didn't work, it used date. Then tried rtcds. (repeat) The time was not synched at all wrt the time zones and also the time. There were 1~3 sec offset besides the TZ problem.

 

  15793   Wed Feb 3 16:27:19 2021 AnchalSummaryBHDSatellite Amplifier Transfer Functions and noise After modifications

I have made modifications recommended in this doc. The changes made are:

  • R24: 19.6k to 4.99k Ohms
  • R20: 19.6k to 4.99k Ohms
  • R23: 787 to 499 Ohms
  • Removed C16.

I took transfer function measurements, fitted them with zeros and poles and plotted it against the zero model of the circuit. The zeros and poles we intended to shift are matching well with 3Hz zero and 30 Hz pole. The later pole at 1500 Hz is at a higher value from what is predicted by zero.

I also took noise measurements and they are in good agreement with the noise predicted by zero.


Edit Wed Feb 10 15:14:13 2021 :

THE NOISE MEASUREMENT WAS WRONG HERE. SEE 40m/15799.

Attachment 1: D1002818_S2100029_TFAfterChanges.pdf
D1002818_S2100029_TFAfterChanges.pdf D1002818_S2100029_TFAfterChanges.pdf D1002818_S2100029_TFAfterChanges.pdf D1002818_S2100029_TFAfterChanges.pdf D1002818_S2100029_TFAfterChanges.pdf D1002818_S2100029_TFAfterChanges.pdf D1002818_S2100029_TFAfterChanges.pdf D1002818_S2100029_TFAfterChanges.pdf
Attachment 2: D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf
D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf
Attachment 3: D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf
D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf
Attachment 4: D1002812_S2100029_After_Modifications_Feb3.jpg
D1002812_S2100029_After_Modifications_Feb3.jpg
Attachment 5: AfterChanges.zip
  15792   Wed Feb 3 15:24:52 2021 gautamUpdateCDSCDS crash and CDS/IFO recovery

Didn't get a chance to comment during the meeting - This was almost certainly a coincidence. I have never had to do this - I assert, based on the ~10 labwide reboots I have had to do in the last two years, that whether the timing errors persist on reboot or not is not deterministic. But this is beyond my level of CDS knowledge and so I'm happy for Rolf / Jamie to comment. I use the reboot script - if that doesn't work, I use it again until the systems come back without any errors.

Quote:

This looked like the usual timing issue. It looked like "ntpdate" is not available in the new system. (When was it updated?)

The hardware clock (RTC) of these hosts are set to be PST while the functional end host showed UTC. So I copied the time of the UTC time from the end to the vertex machines.
For the time adjustment, the standard "date" command was used

> sudo date -s "2021-02-03 07:11:30"

This made the trick. Once IOP was restarted, the "DC" indicators returned to **Green**, restarting the other processes were straight forward and now the CDS indicators are all green.

I don't think this is a problem, the NTP synchronization is handled by timesyncd now.

Quote:

NTP synchronization is not active. Is this OK?

I defer restoring the LSC settings etc since I guess there is not expected to be any interferometer activity for a while.

  15791   Tue Feb 2 23:29:35 2021 KojiUpdateCDSCDS crash and CDS/IFO recovery

I worked around the racks and the feedthru flanges this afternoon and evening. This inevitably crashed c1lsc real-time process.
Rebooting c1lsc caused multiple crashes (as usual) and I had to hard reboot c1lsc/c1sus/c1ioo
This made the "DC" indicator of the IOPs for these hosts **RED**.

This looked like the usual timing issue. It looked like "ntpdate" is not available in the new system. (When was it updated?)

The hardware clock (RTC) of these hosts are set to be PST while the functional end host showed UTC. So I copied the time of the UTC time from the end to the vertex machines.
For the time adjustment, the standard "date" command was used

> sudo date -s "2021-02-03 07:11:30"

This made the trick. Once IOP was restarted, the "DC" indicators returned to **Green**, restarting the other processes were straight forward and now the CDS indicators are all green.

controls@c1iscex:~ 0$ timedatectl
      Local time: Wed 2021-02-03 07:35:12 UTC
  Universal time: Wed 2021-02-03 07:35:12 UTC
        RTC time: Wed 2021-02-03 07:35:26
       Time zone: Etc/UTC (UTC, +0000)
     NTP enabled: yes
NTP synchronized: no
 RTC in local TZ: no
      DST active: n/a

NTP synchronization is not active. Is this OK?


With the recovered CDS, the IMC was immediately locked and the autolocker started to function after a few pokes (like manually running of the "mcup" script). However, I didn't see any light on the AS/REF cameras as well as the test mass faces. I'm sure the IMC alignment is OK. This means the TTs are not well aligned.

So, burtrestored c1assepics with 12:19 snapshot. This immediately brought the spots on the REFL/AS.

Then the arm were aligned, locked, and ASSed. I tried to lock the FP arms. The transmissions were at the level of 0.1~0.3. So some manual alignment of ITMY and BS were necessary. After having the TRs of ~0.8, I still could not lock the arms. The signs of the servo gains were flipped to -0.143 for X arm and -0.012 for Y arm, and the arms were locked. ASS worked well and the ASS offsets were offloaded to the SUSs.

 

  15790   Tue Feb 2 18:24:54 2021 KojiUpdateBHDSOS assembly

You can remove the components of the optical table enclosure (black ones) and use the optical table as your working area too.

 

  15788   Tue Feb 2 17:09:17 2021 yehonathanUpdateBHDSOS assembly

I set up a working area on the table next to the south flow bench (see attachment). I also brought in a rolling table for some extra space.

I covered all the working surfaces with a foil from the big roll between 1x3 and 1x4.

I took the SOSs, SOS parts and the OSEMS from the MC2 table to the working area.

I cleaned some LN Allen keys with isopropanol and put them on the working table, please don't take them.

Attachment 1: 20210202_165501.jpg
20210202_165501.jpg
Attachment 2: 20210202_162452.jpg
20210202_162452.jpg
  15787   Tue Feb 2 11:57:46 2021 AnchalSummaryBHDHAM-A Coil Driver measurements After modifications TF and Noise S2100028

I have made the modifications on the other board D1100687 S2100028 as well. The measurements were taken as mentioned in 40m/15784. All conclusions remain the same as 40m/15784. The attached zip file contains all measurement data, before and after the modifications.


Edit Wed Feb 3 16:44:51 2021 :

Added zero modeled noise in the noise spectrum curves. The acquisition mode curves are in agreement with the model. The noise in Run mode is weirdly lower than predicted by zero.

Attachment 1: D1100687_S2100028_After_Modifications_Feb01_2021.jpg
D1100687_S2100028_After_Modifications_Feb01_2021.jpg
Attachment 2: D1100117_S2100028_TF.pdf
D1100117_S2100028_TF.pdf D1100117_S2100028_TF.pdf D1100117_S2100028_TF.pdf D1100117_S2100028_TF.pdf D1100117_S2100028_TF.pdf
Attachment 3: D1100117_S2100028_Voltage_Noise_Spectrum.pdf
D1100117_S2100028_Voltage_Noise_Spectrum.pdf D1100117_S2100028_Voltage_Noise_Spectrum.pdf D1100117_S2100028_Voltage_Noise_Spectrum.pdf D1100117_S2100028_Voltage_Noise_Spectrum.pdf D1100117_S2100028_Voltage_Noise_Spectrum.pdf
Attachment 4: D1100117_S2100028_Current_Noise_Spectrum.pdf
D1100117_S2100028_Current_Noise_Spectrum.pdf D1100117_S2100028_Current_Noise_Spectrum.pdf D1100117_S2100028_Current_Noise_Spectrum.pdf D1100117_S2100028_Current_Noise_Spectrum.pdf D1100117_S2100028_Current_Noise_Spectrum.pdf
Attachment 5: AfterChanges.zip
  15786   Mon Feb 1 12:30:21 2021 gautamUpdateElectronicsMore careful characterization

Summary:

  1. Swapping out the KEPCO HV supplies (linear) I was using for a pair of HP6209s I borrowed from Rich has improved the noise performance somewhat.
  2. However, there is still an excess relative to the model. I confirmed that this excess originates from the PA95 part of the circuit (see details).
  3. The bypass capacitors don't seem to have any effect on the measured ripple from these HP6209s. Maybe they're internally fitted with some 10uF or similar bypass caps?
  4. The production version of this board, with several improvements (after discussions with Koji and Rich), are on the DCC. They're being fabbed right now and will arrive in ~1 week for more bench testing. 

Power supply bypassing [updated 10pm]:

As mentioned earlier in this thread, I prepared a box with two 10uF, 1kV rated capacitors to bypass the high-voltage rails (see inset in the plot), to see if that improves the performance. However, in measuring the voltage ripple directly with the SR785 (no load connected), I don't see any significant difference whether the decoupling caps are connected or not, see Attachment #1. For this, and all other HV measurements made, I used this box to protect the SR785. One hypothesis is that this box itself is somehow introducting the excess noise, maybe because of leakage currents of the diode pair going into the 1Mohm SR785 input impedance, but I can't find any spec for this, and anyway, these diodes should be at ground potential once the transient has settled and the DC blocking capacitor has charged to its final value.

Note that the 10uF caps have an ESR of 7.2 mOhms. The HP6209 has a source impedance "<20mOhm" when operated as a CV source, per the datasheet. So perhaps this isn't so surprising? The same datasheet suggests the source impedance is 500 mOhms from 1kHz to 100 kHz, so we should see some improvement there, but I only measured out to 2 kHz, and I didn't take much effort to reduce these crazy peaks so maybe they are polluting the measurement out there. There must also be some continuous change of impedance, it cannot be <20 mOhm until 1 kHz and then suddenly increase to 500 mOhms. Anyways, for this particular circuit, the nosie DC-1kHz is what is important so I don't see a need to beat this horse more. 

Simplified circuit testing:

I decided to see if I can recover the spec'd voltage noise curve from the PA95 datasheet. For this, I configured the PA95 as a simple G=31 non-inverting amplifier (by not stuffing the 15 uF capacitor in the feedback path). Then, with the input grounded, I measured the output voltage noise on the circuit side of the 25kohm resistor (see inset in Attachment #2). To be consistent, I used the DC blocking box for this measurement as well, even though the output of the PA95 under these test conditions is 0V. Once again, there is considerable excess around ~100 Hz relative to a SPICE model. On the basis of this test, I think it is fair to say that the problem is with the PA95 itself. As far as I can tell, I am doing everything by the book, in terms of having gain > 10, using a sufficiently large compensaiton cap, HV rail decoupling etc etc. Note that the PA95 is a FET input opamp, so the effects of input current noise should be negligible. The datasheet doesn't provide the frequency dependence, but if this is just shot noise of the 1200 pA input bias current (for 300 V rails, per the spec), this is totally negligible, as confirmed by LTspice.

In the spirit of going step-by-step, I then added the feedback capacitor, and still, measured noise in excess of what I would expect from my model + SR785 measurement noise.

Integrated circuit testing:

After the above simplified test, I stuffed a full channel as designed, and tested the noise for various drive currents. To best simulate the operating conditions, an Acromag XT1541 was used to set the DC voltage that determines the drive current through the 25 kohm resistor. The measurements were made on the circuit side of this resistor (I connected a 20ohm resistor to ground to simulate the OSEM). As shown in Attachment #3, the noise with these HP6209 supplies is significantly better than what I saw with the KEPCO supplies, lending further credence to the hypothesis that insufficient PSRR is the root of the problem here. I've added subplots in a few different units - to be honest, I think that reaching this level of measured displacement noise at the 40m at 100 Hz would already be pretty impressive.

So what's next?

The main design change is that a passive R-C-R (4k-3uF-20k) replaces the single 25kohm resistor at the output of the PA95. 

  • This allows similar current drive range.
  • But adds an LPF to filter out the observerd excess noise at 100 Hz. 

Let's see if this fixes the issue. Not that I've also added a pair of input protection diodes to the input of the PA95 in the new design. The idea is that this would protect the (expensive) PA95 IC from, for example, the unit being powered with the +/- 18V rail but not the +/- 300 V rail. As I type this, however, I wonder if the leakage current noise of these diodes would be a problem. Once again, the datasheet doesn't provide any frequency dependence, but if it's just the shot noise of the 1nA expected when the diodes are not reverse biased (which is the case when the PA95 is operating normally since both inputs are at nearly the same potential), the level is ~20 fA/rtHz, comparable to the input current noise of the PA95, so not expected to be an issue. In the worst case, the PCB layout allows for this component to just be omitted. 

Attachment 1: HVPS.pdf
HVPS.pdf
Attachment 2: HV_testckt.pdf
HV_testckt.pdf
Attachment 3: totalNoise.pdf
totalNoise.pdf
  15785   Fri Jan 29 17:57:17 2021 AnchalHowToCDSAcromag wiring investigation

I found a white paper  from Acromag which discusses how to read differential signal using Acromag units. The document categorically says that differential signals are always supposed to be transmitted in three wires. I provides the two options of either using the RTN to connect to the signal ground (as done in Attachment 3) or locally place 10k-100k resistors between return and IN+ and IN- both (Attachment 2).

I have provided possible scenarios for these.

Using two wires to carry differential signal (Attachment 1):

  • I assume this is our preferential way to connect.
  • We can also assume all single-ended inputs as differential and do a signal condition agnostic wiring.
  • Attachment 3 show what were the results for different values of resistors when a 2Hz 0.5V amplitude signal from SR785 which as converted to differential signal using D1900068 was measured by acromag.
  • The connection to RTN is symmetrical for both inputs.

Using three wires to carry differential signal (Attachment 2):

  • This is recommended method by the document in which it asks to carry the GND from signal source and connect it to RTN.
  • If we use this, we'll have to be very cautious on what GND has been shorted through the acromag RTN terminals.
  • This would probably create a lot of opportunities for ground loops to form.

Using an acromag card without making any connection with RTN is basically not allowed as per this document.

Attachment 1: GeneralLabWiringDiffWith2Wires.pdf
GeneralLabWiringDiffWith2Wires.pdf
Attachment 2: GeneralLabWiringDiffWith3Wires.pdf
GeneralLabWiringDiffWith3Wires.pdf
Attachment 3: DiffReadResistorbtwnINandRTN.pdf
DiffReadResistorbtwnINandRTN.pdf DiffReadResistorbtwnINandRTN.pdf DiffReadResistorbtwnINandRTN.pdf
  15784   Fri Jan 29 15:39:30 2021 AnchalSummaryBHDHAM-A Coil Driver measurements After modifications TF and Noise S2100027

I fitted zeros and poles in the measured transfer function of D1100687 S2100027 and got zeros at 130 Hz and 234 Hz and poles at 10Hz and 2845 Hz. These values are different from the aimed values in this doc, particularly the 234Hz zero which was aimed at 530 Hz in the doc.

I also took the noise measurement using the same method as described in 40m/15780. The noise in Acquisition mode seems to have gone up in 10 Hz - 500 Hz region compared to the measurement in 40m/15780 before the modifications.

All channels are consistent with each other.


Edit Mon Feb 1 12:24:14 2021:
Added zero model prediction after the changes. The measurements match with the predictions.


Edit Wed Feb 3 16:46:59 2021:

Added zero modeled noise in the noise spectrum curves. The acquisition mode curves are in agreement with the model. The noise in Run mode is weirdly lower than predicted by zero.

Attachment 1: D1100687_S2100027_After_Modifications_Jan28.jpg
D1100687_S2100027_After_Modifications_Jan28.jpg
Attachment 2: D1100117_S2100027_TF.pdf
D1100117_S2100027_TF.pdf D1100117_S2100027_TF.pdf D1100117_S2100027_TF.pdf D1100117_S2100027_TF.pdf D1100117_S2100027_TF.pdf
Attachment 3: D1100117_S2100027_Voltage_Noise_Spectrum.pdf
D1100117_S2100027_Voltage_Noise_Spectrum.pdf D1100117_S2100027_Voltage_Noise_Spectrum.pdf D1100117_S2100027_Voltage_Noise_Spectrum.pdf D1100117_S2100027_Voltage_Noise_Spectrum.pdf D1100117_S2100027_Voltage_Noise_Spectrum.pdf
Attachment 4: D1100117_S2100027_Current_Noise_Spectrum.pdf
D1100117_S2100027_Current_Noise_Spectrum.pdf D1100117_S2100027_Current_Noise_Spectrum.pdf D1100117_S2100027_Current_Noise_Spectrum.pdf D1100117_S2100027_Current_Noise_Spectrum.pdf D1100117_S2100027_Current_Noise_Spectrum.pdf
Attachment 5: AfterChanges.zip
  15783   Thu Jan 28 22:34:21 2021 gautamUpdateSUSDe-whitening

Summary:

  1. We will need de-whitening filters for the BHD relay optics in order to meet the displacement noise requirements set out in the DRD. I think these need not be remotely switchable (depends on specifics of LO phase control scheme). SR2, PR2 and PR3 can also have the same config, and probably MC1, MC3 as well.
  2. We will need de-whitening filters for the non test mass core IFO optics (PRM, SRM, BS, and probably MC2).
  3. I am pretty sure we will not be able to have sufficient DAC range for the latter class of optics if we have to:
    1. Supply the DC bias.
    2. Do the LSC and ASC actuation in the presence of reasonable sensing noise levels.
    3. Engage de-whitening to low-pass-filter the DAC noise at ~200 Hz.

Details:

Attachment #1 shows the DAC noise models for the General Standards 16-bit and 18-bit DACs we are expecting to have.

  • The 16-bit model has been validated by me at the 40m a few years ago.
  • We have never used the 18-bit flavor at the 40m, and there are all manner of quirks apparently related to zero crossings and such. So the noise may be up to x2 higher (we won't have as much freedom necessarily as the sites to bias the DAC on one side of the zero crossing if we also need to use the same DAC channel to supply the DC bias current for alignment.

Attachment #2 shows the expected actuation range for DC optic alignment, assuming we use the entire DAC range for this purpose.

  • Clearly, we need to do other things with the same DAC channels as well, so this is very much an upper bound of what will be possible.
  • Let's assume we will not go lower than 100ohms.
  • For all new optics we are suspending, we should aim to get the pitch balancing to within 500urad. With a 2x2m=4m optical lever arm, this corresponds to a 2mm spot shift. Should be doable.
  • This could turn out to be a serious problem for PRM, BS and SRM if we hope to measure squeezing - the <AUX DOF>-->DARM coupling could be at the level of -40dB, and at 200 Hz, the DAC noise would result in PRCL/MICH/SRCL noise at the level of ~10^-15m/rtHz, which would be 10^-17m/rtHz in DARM. I don't think we can get 20dB of feedforward cancellation at these frequencies. For demonstrating locking using a BHD error signal, maybe this is not a big deal.

Attachment #3 shows the current and proposed (by me, just a rough first pass, not optimized in any way yet) de-whitening filter shapes. These shapes can be tweaked for sure.

  • The existing de-whitening filter is way too aggressive. FWIW, the DRD "models" a "4th order Chebyshev low pass filter" which doesn't exist anywhere as far as I know.
  • Since the DAC noise is below 1 uV/rtHz at all frequencies of interest, we never need to have >60dB de-whitening anywhere as the input referred noise of any circuit we build will exceed 1 nV/rtHz.
  • I propose 3 poles, 3 zeros. In the plot, these poles are located at 30Hz, 50Hz, 2kHz, and the zeros are at 300 Hz, 300 Hz, 800 Hz. 
  • The de-whitening is less agressive below 100 Hz, where we still need significant LSC actuation ability. Considering the sensing noise levels at the 40m, I don't know if we can have reasonable LSC and ASC loop shapes and still have the de-whitening.
  • Once again, PRM, SRM and BS will be the most challenging.
  • For the BHD relay optics, once we have the de-whitening, we won't have the option of turning on a high-frequency (~kHz) dither line because of insufficient DAC range. 

Attachment #4 puts everything into displacement noise units. The electronics noise of the coil driver / de-whitening circuit have not been included so at high frequencies, the projection is better than what will actually be realizable, but still well below the BHD requirement of 3e-17 m/rtHz.

Attachment 1: DACnoiseModels.pdf
DACnoiseModels.pdf
Attachment 2: actuationRange.pdf
actuationRange.pdf
Attachment 3: deWhiteTFs.pdf
deWhiteTFs.pdf
Attachment 4: dispNoiseModels.pdf
dispNoiseModels.pdf
  15782   Thu Jan 28 21:44:45 2021 gautamSummaryBHDHAM-A Coil Driver measurements After modifications

Looks fine to me visually but the verdict can only be made once the z:p locations are quantitatively confirmed, and the noise tests pass. It would be interesting to see what kind of time-domain transient (in N of force) switching on the de-whitening introduces, i guess best done interferometrically.

Quote:

I'll wait for comments until tomorrow to proceed with changes in the other board as well. I'll do noise measurements tomorrow.

  15781   Thu Jan 28 18:04:55 2021 AnchalSummaryBHDHAM-A Coil Driver measurements After modifications

I did the recommended modifications on of the boards with serial number S2100028. These included:

  • R13, R27: 160 -> 75
  • C11, C21: 470 nF -> 68nF
  • C19: 4.7 uF -> 470 nF
  • R15: 3.23 kOhm -> 1.82 kOhm

I took transfer function measurements with same method as in 40m/15774 and I'm presenting it here to ensure the modifications are correct and if I should proceed to the next board as well. I didn't have the data used to make plots in here but I think the poles and zeros have landed in the right spot. I'll wait for comments until tomorrow to proceed with changes in the other board as well. I'll do noise measurements tomorrow.

Attachment 1: D1100117_S2100027_TF.pdf
D1100117_S2100027_TF.pdf D1100117_S2100027_TF.pdf D1100117_S2100027_TF.pdf D1100117_S2100027_TF.pdf D1100117_S2100027_TF.pdf
Attachment 2: AfterChanges.zip
  15780   Thu Jan 28 12:53:14 2021 AnchalSummaryBHDHAM-A Coil Driver measurements before modifications

I took some steps to reduce the coupling of 60 Hz harmonics in noise measurement. The box was transferred to the floor instead of on top of another instrument. Measurement was immediately converted into single-ended using SR560 in battery mode with a gain of 10. All of the setups was covered in aluminum foil to increase isolation.

Spectrum measurement details

 

Attachment 1: D1100117_S2100027_Current_Noise_Spectrum.pdf
D1100117_S2100027_Current_Noise_Spectrum.pdf D1100117_S2100027_Current_Noise_Spectrum.pdf D1100117_S2100027_Current_Noise_Spectrum.pdf D1100117_S2100027_Current_Noise_Spectrum.pdf D1100117_S2100027_Current_Noise_Spectrum.pdf
Attachment 2: D1100117_S2100027_Voltage_Noise_Spectrum.pdf
D1100117_S2100027_Voltage_Noise_Spectrum.pdf D1100117_S2100027_Voltage_Noise_Spectrum.pdf D1100117_S2100027_Voltage_Noise_Spectrum.pdf D1100117_S2100027_Voltage_Noise_Spectrum.pdf D1100117_S2100027_Voltage_Noise_Spectrum.pdf
Attachment 3: D1100117_S2100028_Current_Noise_Spectrum.pdf
D1100117_S2100028_Current_Noise_Spectrum.pdf D1100117_S2100028_Current_Noise_Spectrum.pdf D1100117_S2100028_Current_Noise_Spectrum.pdf D1100117_S2100028_Current_Noise_Spectrum.pdf D1100117_S2100028_Current_Noise_Spectrum.pdf
Attachment 4: D1100117_S2100028_Voltage_Noise_Spectrum.pdf
D1100117_S2100028_Voltage_Noise_Spectrum.pdf D1100117_S2100028_Voltage_Noise_Spectrum.pdf D1100117_S2100028_Voltage_Noise_Spectrum.pdf D1100117_S2100028_Voltage_Noise_Spectrum.pdf D1100117_S2100028_Voltage_Noise_Spectrum.pdf
Attachment 5: SpectrumMeasurement.zip
  15779   Tue Jan 26 15:37:25 2021 AnchalHowToCDSAcromag wiring investigation

Here I present few wiring diagrams when using Acromag to avoid noisy behavior and ground loops.


Case 1: Only single-ended sources

  • Attachment 1 gives a functioning wiring diagram when all sources are single ended.
  • One should always short the RTN to IN- pin if the particular GND carried by that signal has not been shorted before to RTN for some other signal.
  • So care is required to mark different GNDs of different powersupply separately and follow where they inadvertently get shorted, for example when a photodiode output is connected to FSS Box.
  • Acromag should serve as the node of all GNDs concerned and all these grounds must not be connected to Earth GND at power supply ends or in any of the signal sources.
  • I think this is a bit complicated thing to do.

Case 2: Some single and some differential sources

  • Connect all single ended sources same as above keeping care of not building any ground loops.
  • The differential source can be connected to IN+ and IN- pins, but there should be a high resistance path between IN- and RTN. See Attachment 2.
  • Why this is the case, I'm not sure since I could not get access to internal wiring of Acromag (no response from them). But I have empirically found this.
  • This helps IN- to float with respect to RTN according to the negative signal value. I've found that 10kOhm resistance works good. See 40m/15778.
  • If RTN is shorted to nearby Earth GND (assuming none of the other power supply GNDs have been shorted to Earth GND) shows a reduction in noise for differential input. So this is recommended.
  • This wiring diagram carries all complexity of previous case along with the fact that RTN and anything connected to it is at Earth GND now.

Case 3: Signal agnostic wiring

  • Attachment 3 gives a wiring diagram which mimics the high resistance shorting of RTN to IN- in all ports regardless of the kind of signal it is used for reading.
  • In this case, instead of being the node of the star configuration for GND, acromag gets detached from any ground loops.
  • All differences in various GNDs would be kept by the power supplies driving small amounts of current through the 10 kOhm resistors.
  • This is a much simpler wiring diagram as it avoids shorting various signal sources or their GNDs with each other, avoiding some of the ground loops.

Edit Wed Jan 27 13:38:19 2021 :

This solution is not acceptable as well. Even if it is successfull in reading the value, connecting resistor between IN- and RTN will not break the ground loops and the issue of ground loops will persist. Further, IN- connection to RTN breaks the symmetry between IN-  and IN+, and hence reduces the common mode rejection which is the intended purpose of differential signal anyways. I'll work more on this to find a way to read differential signals without connecitng IN- and RTN. My first guess is that it would need the GND on the source end to be connected to EarthGND and RTN on acromag end to be connected to EarthGND as well.

 

Attachment 1: GeneralLabWiring.pdf
GeneralLabWiring.pdf
Attachment 2: GeneralLabWiring2.pdf
GeneralLabWiring2.pdf
Attachment 3: GeneralLabWiring3.pdf
GeneralLabWiring3.pdf
  15778   Tue Jan 26 12:59:51 2021 AnchalHowToCDSAcromag wiring investigation

Taking inspiration from SR785 on how it reads differential signal, I figured that acromag too always need a way to return current through RTN ports always. That must be the reason why everything goes haywire when RTN is not connected to IN-. Now for single ended signals, we can always short RTN to IN- and keep same GND but then we need to be careful in avoiding ground loops. I'm gonna post a wiring diagram in next post to show how if two signal sources connect to each other separately, a GND loop can be formed if we tie each IN- port to RTN on an acromag.
Coming to the issue of reading a differential signal, what SR785 does is that it connects 50 Ohm resistance between Earth GND and differential signal shields (which are supposed to signal GND). In a floating GND setting, SR785 connects a 1 MOhm resistor between input shield and Earth GND. This can be used to read a differential signal through a single BNC cable since the shiled can take arbitrary voltages thanks ti the 1 MOhm resistor.

We can do the same in acromag. Instead of shorting RTN to IN- ports, we can connect them through a large resistor which would let IN- float but will give a path for current to return through RTN ports. Attached here are few scenarios where I connected IN- to RTN throguh wire, 820 Ohms, 10kOhms and 1MOhms in two sub cases where RTN was left open or was shorted to Earth GND. In all cases, the signal was produced by a 9V battery outputing roughly 8.16V. It seems that 10kOhm resistor between RTN and IN- with RTN connected to Earth GND is the best scenario noise wise. I'll post more results and a wiring diagram soon.

Attachment 1: TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf
TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf
  15777   Tue Jan 26 10:58:30 2021 gautamUpdateSUSMC2 tickler stuck on

For whatever reason, the autolocker didn't turn the tickle off for several hours. Seems to work okay now. The linked plot suggests that the coil balancing on MC2 is pretty lousy.

  15776   Mon Jan 25 18:18:04 2021 AnchalSummaryBHDSatellite Amplifier Transfer Functions and noise

 

I took transfer function and noise measurement of satellite amplifier box's photodiode transimpedance circuit. For the measurement, I created a makeshift connector to convert backside DB25 into DB9 with the 4 channels for PDA input. The output was taken in differential form at the front PD Output port. To feed current to the circuit, I put in 12 kOhm resistors in series at the inputs, so the V/V transfer function measured was multiplied by 12 kOhm to get the transimpedance of the circuit.


Transfer Function Measurement details

  • SR785 source out was fed into PDA input pins using a makeshift BNC-DB9-DB25 converter.
  • The output from PDOut DB9 port was fed to test switch in D1900068 to separate differential signal.
  • This differential signal was fed back to SR785 at input 2 in A-B configuration.
  • Measurements are taken with file D1002818_TF.yml and D1002818_TF_LF.yml.
  • A measurement of just cables without the DUT is taken as well.
  • Commands.txt list all the commands used.
  • All data is compiled and plotted in Plotting.ipynb
  • D1100117_S2100029_TFandNoiseSpectrum.pdf shows all the transfer functions measured.

Spectrum Measurements

  • Two pair of BNC cables were twisted together and clips were added at ends.
  • One of the GND was connected to board GND. Rest were left unconnected to avoid ground loops.
  • Each pair of signal was connected to PDOutP/N.
  • The PDA inputs were shorted together to make zero input current to the board.
  • Instrument noise with cables was measured by shorting the clips of the center cores and one of the shields of the two BNC cables together.
  • Measurements were taken with file D1002818_SP.yml and D1002818_SP_LF.yml.
  • Input referred current noise spectrum was calculated by dividing the output voltage noise spectrum by the measured transfer function.
  • D1100117_S2100029_TFandNoiseSpectrum.pdf shows all the output votlage noise spectrum and input referred current noise spectrum measured.

Edit Wed Feb 10 15:14:13 2021 :

THE NOISE MEASUREMENT WAS WRONG HERE. SEE 40m/15799.

Attachment 1: D1002818_S2100029_TFandNoiseSpectrum.pdf
D1002818_S2100029_TFandNoiseSpectrum.pdf D1002818_S2100029_TFandNoiseSpectrum.pdf D1002818_S2100029_TFandNoiseSpectrum.pdf
Attachment 2: D1002818_Testing.zip
  15775   Wed Jan 20 19:12:16 2021 gautamUpdateCDSSLwebview fixed

After fixing multiple issues, the model webviews are updating, should be done by tomorrow. It should be obvious from the timestamps on the index page which are the new ones. These new screens are better than the old ones and offer more info/details. Please look at them, and let me know if there are broken links etc. Once we are happy with this new webview, we can archive the old files and clean up the top directory a bit. I don't think this adds anything to the channel accounting effort but it's a nice thing to have up-to-date webviews, I found the LLO ones really useful in setting up the AS WFS model.

BTW, the crontab on megatron is set to run every day at 0844. The process of updating the models is pretty heavy because of whatever MATLAB overhead. Do we really need to have this run every day? I modified the crontab to run every other Saturday, and we can manually run the update when we modify a model. Considering this hasn't been working for ~3 years, I think this is fine, but if anyone has strong preference you can edit the crontab.

If someboy can volunteer to fix the MEDM screenshot that would be useful.

ELOG V3.1.3-